# Optuna vs Code Llama: Are LLMs a New Paradigm for Hyperparameter Tuning?

### v3

@@ 4.3 Model Performance Analysis
Models from train dataset. The comparison between models included in fine-tuning along with their 5-epoch versions (AlexNet, RNN, and ConvNeXt) and models whose 5-epoch versions were excluded from fine-tuning (ResNet, RegNet, MobileNetV2) highlights the impact of fine-tuning on RMSE values. As shown in Figure 3(a), models that underwent fine-tuning demonstrate consistently lower RMSE values, with significantly reduced confidence intervals, reflecting improved stability and robustness in hyperparameter recommendations. ConvNeXt, in particular, exhibits a notable decrease in RMSE, confirming the effectiveness of fine-tuning. Models whose 5-epoch versions were not used in fine-tuning (Figure 3(b)) exhibit higher RMSE values in the LLM’s predictions compared to Optuna, indicating room for improvement in generalization. Nonetheless, the fine-tuned model maintains reasonable stability, as reflected in confidence intervals that, while broader, still reflect consistent behavior across trials.
@@
Models from test dataset. To evaluate the performance of the fine-tuned model, we tested it on three previously unseen architectures: InceptionV3, VisionTransformer, and MaxViT. These models were not part of the fine-tuning process, meaning Code Llama had no prior exposure to them. The fine-tuned model generated corresponding hyperparameters, and we compared the resulting RMSE values against those obtained using the Optuna framework.
As shown in Figure 3(c), the most notable improvement is observed for VisionTransformer and MaxViT, where fine-tuned Code Llama achieves significantly lower RMSE values than Optuna, indicating that its hyperparameter recommendations lead to more stable and effective optimization for these architectures. Conversely, for InceptionV3, Optuna demonstrates a clear advantage, yielding a lower RMSE, although with wider confidence intervals, suggesting greater variability in its predictions. This contrast highlights that while fine-tuned Code Llama successfully generalizes to certain unseen models, its effectiveness may depend on the architectural similarity to those encountered during training. Nevertheless, its ability to generate competitive hyperparameter recommendations without additional tuning reinforces its potential for efficient and automated hyperparameter selection, particularly in scenarios requiring rapid model deployment.
@@ 4.4 Prediction Dynamics Across Epochs
In this experiment, we analyze hyperparameter recommendations for a diverse set of models. While the full evaluation was conducted across all models in our study, we present detailed results for four representative architectures: MobileNetV2 and DenseNet, which were part of the training set (participated in Code Llama fine-tuning with data from 1 and 2 epochs), and MaxVit and VisionTransformer, which belong to the test set (not involved in Code Llama fine-tuning). This selection provides a balanced comparison between models seen and unseen during fine-tuning.
The experiment illustrates the accuracy of one-shot predictions across multiple epochs (1, 2, 5, 10, 15, and 20), as obtained using the fine-tuned Code Llama, and compares them to the best and average accuracy values achieved by Optuna. As shown in Figure 4, training models (MobileNetV2 and DenseNet) exhibit a consistent increase in prediction accuracy as the number of epochs increases. Their performance surpasses the average Optuna accuracy and closely approaches the best Optuna results, confirming the effectiveness of fine-tuning on seen models.
For test models (MaxVit and VisionTransformer), Code Llama also demonstrates strong generalization. Notably, MaxVit achieves one-shot accuracy close to 0.7, significantly outperforming both the average and best Optuna results, which remain at 0.1. VisionTransformer also steadily improves with more training epochs, showing superior performance relative to Optuna’s average and best scores. Detailed results for additional models can be found in the Supplementary Material (Figures 67). These outcomes highlight Code Llama’s ability to generate competitive hyperparameter configurations even for unseen architectures.
While exhaustive search can theoretically yield optimal hyperparameter values, the associated time costs are a considerable constraint. In real-world model training, exhaustive search becomes highly impractical, underscoring the value of efficient hyperparameter tuning methods like Code Llama.
@@ 4.5 Training Trends and Performance Comparison
As part of the experiment, we decided to compare the results of 100 training runs obtained with Optuna with the best values of 100 hyperparameter predictions after fine-tuning with Code Llama, as well as the one-shot prediction accuracy of fine-tuned Code Llama. One-shot metrics were also obtained for each model, which provide an indication of how the models can perform when predicted one time using the hyperparameters provided by Code Llama.
@@
Refer to caption Refer to caption
Figure 5: Comparative analysis of the accuracy of 6 computer vision models using Optuna (green) on 100 trials, the best accuracy value on hyperparameters obtained from fine-tuned Code Llama ( blue) and a one-shot prediction (purple) based on fine-tuned Code Llama.
@@
Optuna, as seen from the green dots in the graphs, shows a significant spread in accuracy values, especially for a model like ConvNeXt and MobileNetV3, indicating high variability in results. At the same time, fine-tuning shows more stable results, confirming its effectiveness in achieving consistent accuracy. The greatest effect of fine-tuning is seen for DenseNet and MobileNetV3, where the blue line is noticeably higher than the purple and green ones, indicating significant improvements. For some models, like GoogLeNet, the difference between the best values after fine-tuning and the one-shot predictions is less pronounced, indicating limited potential for improvement for individual architectures, but even here fine-tuning maintains a small advantage in accuracy. Detailed results for additional models can be found in the Supplementary Material (Figures 8 - 10).
@@ 5 Conclusion
Our findings demonstrate that the fine-tuned Code Llama is highly effective for identifying optimal hyperparameter values across diverse machine learning models, achieving improvements in both accuracy and stability. The fine-tuned model not only enhances training efficiency but also delivers competitive performance across various scenarios, making it a valuable asset for researchers and practitioners.
Our evaluations reveal that fine-tuned Code Llama often meets or exceeds the accuracy achieved by Optuna, a well-established hyperparameter optimization framework. Notably, fine-tuning cycles resulted in higher accuracy and more stable predictions across multiple models, including complex architectures like ResNet and DenseNet, where fine-tuning significantly surpassed baseline results. Additionally, one-shot predictions using Code Llama demonstrated accuracy levels close to those obtained through full fine-tuning, highlighting its potential for rapid and effective optimization.
A key insight from this study is the efficiency advantage of fine-tuned Code Llama. While Optuna typically requires training across 100 trials to identify suitable hyperparameters, our approach achieves comparable or superior results with significantly fewer computational resources. This efficiency enables faster decision-making and resource allocation, particularly in time-sensitive or resource-constrained environments.
Overall, our study underscores the effectiveness of fine-tuned Code Llama as a powerful tool for hyperparameter optimization. Its ability to provide high-quality, stable results with minimal training iterations positions it as a compelling alternative to traditional optimization methods, advancing machine learning efficiency while making AI more accessible for real-world applications.

### v4

@@ 4.3 Model Performance Analysis
Models from train dataset. The comparison between models included in fine-tuning along with their 5-epoch versions (AlexNet [40], RNN [18], and ConvNeXt [46]) and models whose 5-epoch versions were excluded from fine-tuning (ResNet [27], RegNet [53], MobileNetV2 [57]) highlights the impact of fine-tuning on RMSE values. As shown in Figure 3(a), models that underwent fine-tuning demonstrate consistently lower RMSE values, with significantly reduced confidence intervals, reflecting improved stability and robustness in hyperparameter recommendations. ConvNeXt, in particular, exhibits a notable decrease in RMSE, confirming the effectiveness of fine-tuning. Models whose 5-epoch versions were not used in fine-tuning (see Figure 3(b)) exhibit higher RMSE values in the LLM’s predictions compared to Optuna, indicating room for improvement in generalization. However, the fine-tuned model maintains reasonable stability, as reflected in confidence intervals that, while broader, still reflect consistent behavior across trials.
@@
Models from test dataset. To evaluate the performance of the fine-tuned model, we tested it on three previously unseen architectures: InceptionV3 [62], VisionTransformer [17], and MaxViT [64]. These models were not part of the fine-tuning process, which means that Code Llama had no prior exposure to them. The fine-tuned model generated the corresponding hyperparameters, and we compared the resulting RMSE values against those obtained using the Optuna framework.
As shown in Figure 3(c), the most notable improvement is observed for VisionTransformer and MaxViT, where the fine-tuned Code Llama achieves significantly lower RMSE values than Optuna, indicating that its hyperparameter recommendations lead to more stable and effective optimization for these architectures. Conversely, for InceptionV3, Optuna demonstrates a clear advantage, yielding a lower RMSE, although with wider confidence intervals, suggesting greater variability in its predictions. This contrast highlights that while fine-tuned Code Llama successfully generalizes to certain unseen models, its effectiveness may depend on the architectural similarity to those encountered during training. Nevertheless, its ability to generate competitive hyperparameter recommendations without additional tuning reinforces its potential for efficient and automated hyperparameter selection, particularly in scenarios requiring rapid model deployment.
@@ 4.4 Prediction Dynamics Across Epochs
In this experiment, we analyze hyperparameter recommendations for a diverse set of models. Although a complete evaluation was performed for all models in our study, we present detailed results for four representative architectures: MobileNetV2 and DenseNet [31], which were part of the training set (participated in Code Llama fine-tuning with data from 1 and 2 epochs), and MaxViT and VisionTransformer, which belong to the test set (not involved in Code Llama fine-tuning). This selection provides a balanced comparison between models seen and unseen during fine-tuning.
The experiment illustrates the accuracy of one-shot predictions across multiple epochs (1, 2, 5, 10, 15, and 20), as obtained using the fine-tuned Code Llama, and compares them to the best and average accuracy values achieved by Optuna. As shown in Figure 4, training models (MobileNetV2 and DenseNet) show a consistent increase in prediction accuracy as the number of epochs increases. Their performance surpasses the average Optuna accuracy and closely approaches the best Optuna results, confirming the effectiveness of fine-tuning on seen models.
For test models (MaxVit and VisionTransformer), Code Llama also demonstrates strong generalization. Notably, MaxVit achieves a one-shot accuracy close to 0.7, significantly outperforming both the average and the best Optuna results, which remain at 0.1. VisionTransformer also steadily improves with more training epochs, showing superior performance relative to Optuna’s average and best scores. Detailed results for additional models can be found in the Supplementary Material (Figures 6–7). These outcomes highlight Code Llama’s ability to generate competitive hyperparameter configurations even for unseen architectures.
Although an exhaustive search can theoretically yield optimal hyperparameter values, the associated time costs are a considerable constraint. In real-world model training, exhaustive search becomes highly impractical, underscoring the value of efficient hyperparameter tuning methods such as Code Llama.
@@ 4.5 Training Trends and Performance Comparison
As part of the experiment, we decided to compare the results of 100 training runs obtained with Optuna with the best values of 100 hyperparameter predictions after fine-tuning with Code Llama, as well as the one-shot prediction accuracy of fine-tuned Code Llama. One-shot metrics were also obtained for each model, providing an indication of how the models can perform when predicted once using the hyperparameters provided by Code Llama.
@@
Figure 5: Comparative analysis of the accuracy of 4 computer vision models using Optuna (green) on 100 trials, the best accuracy value on hyperparameters obtained from fine-tuned Code Llama ( blue) and a one-shot prediction (purple) based on fine-tuned Code Llama.
@@
Optuna, as seen from the green dots in the graphs, shows a significant spread in accuracy values, especially for a model such as AlexNet, GoogLeNet, and MobileNetV3, indicating high variability in the results. At the same time, fine-tuning shows more stable results, confirming its effectiveness in achieving consistent accuracy. The greatest effect of fine-tuning is seen for DenseNet and MobileNetV3, where the blue line is noticeably higher than the purple and green ones, indicating significant improvements. For some models, like GoogLeNet, the difference between the best values after fine-tuning and the one-shot predictions is less pronounced, indicating limited potential for improvement for individual architectures, but even here fine-tuning maintains a small advantage in accuracy. Detailed results for additional models can be found in the Supplementary Material (Figures 8-10).
@@ 5 Conclusion
Our findings demonstrate that the fine-tuned Code Llama is highly effective in identifying optimal hyperparameter values across diverse machine learning models, achieving improvements in both accuracy and stability. The fine-tuned model not only enhances training efficiency, but also delivers competitive performance in various scenarios, making it a valuable asset for researchers and practitioners.
Our evaluations reveal that fine-tuned Code Llama often meets or exceeds the accuracy achieved by Optuna, a well-established hyperparameter optimization framework. In particular, fine-tuning cycles resulted in higher accuracy and more stable predictions across multiple models, including complex architectures such as ResNet and DenseNet, where fine-tuning significantly surpassed baseline results. These architectures are commonly used in image restoration and manipulation tasks, highlighting the practical relevance of our method to real-world vision applications.
Additionally, one-shot predictions using Code Llama demonstrated accuracy levels close to those obtained through full fine-tuning, highlighting its potential for rapid and effective optimization.
A key insight from this study is the efficiency advantage of fine-tuned Code Llama. Although Optuna typically requires training across 100 trials to identify suitable hyperparameters, our approach achieves comparable or superior results with significantly fewer computational resources. This makes the method especially attractive for image enhancement workflows, where perceptual quality and resource efficiency must be balanced.
In general, our study underscores the effectiveness of fine-tuned Code Llama as a powerful tool for hyperparameter optimization. Its ability to provide stable high-quality results with minimal training iterations positions it as a compelling alternative to traditional optimization methods, contributing to the development of smarter, faster, and more adaptive pipelines for image manipulation and restoration.
Acknowledgments. This work was partially supported by the Alexander von Humboldt Foundation.